OpenStudy (anonymous):

Find the standard form of the following equation: x^2+y^2+4x+10y+28=0

OpenStudy (anonymous):

Yes, I need to write the standard form and then graph the circle.

OpenStudy (anonymous):

oops i meant \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

first rewrite as \[x^2+4x+y^2+10y=-28\] then complete the square twice do you know how to do that?

OpenStudy (anonymous):

No i don't, I seem to be stuck and not understanding it

OpenStudy (anonymous):

ok i can walk you through it, it is not too hard

OpenStudy (anonymous):

just the mechanics. not the reasoning lets look at \(x^2+4x\) what is half of \(4\) ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

ok so we write \((x+2)^2\) now what is \(2^2\) ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

ok good so when we replace the \(x^2+4x\) on the left with \((x+2)^2\) we have to add \(4\) to the other side

OpenStudy (anonymous):

\[(x+2)^2+y^2+10y=-28+4\] or \[(x+2)^2+y^2+10y=-24\] now we repeat the process with \(y^2+10y\)

OpenStudy (anonymous):

what is half of \(10\) ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

so we are going to replace \(y^2+10y\) by \((y+5)^2\) now that is \(5^2\) ?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

then when we replace \(y^2+10y\) by \((y+5)^2\) we have to add \(25\) to the other side just like before now we have \[(x+2)^2+(y+5)^2=-24+25\] or \[(x+2)^2+(y+5)^2=1\] as we want

OpenStudy (anonymous):

Wow you made it seem so simple, thank you! How do I graph it now?

OpenStudy (anonymous):

it is a circle with center at \((-2,-5)\) and radius \(1\) draw that

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw